# BlockFusion: Expandable 3D Scene Generation using Latent Tri-plane Extrapolation

### 31 Jan 2024 v2

@@ 1. Introduction
Generating large amount of high-quality 3D contents is key for many practical applications, including video-games, film-making, augmented and virtual reality (AR/VR). The increasing demand for high-quality digital contents has made 3D generation a significant topic of research. Recently, in the 2D domain, denoise diffusion models (Sohl-Dickstein et al., 2015), have demonstrated remarkable results in image synthesis and beyond, leading to the development of production-ready 2D generation tools, such as Stable Diffusion (Rombach et al., 2022a), Midjourney, and Dall-E (Ramesh et al., 2021). The success in 2D domain has significantly sparked interest in the development of 3D generation tools. A multitude of researches on 3D generation have been published recently, most notable works include DreamFusion (Poole et al., 2022), Rodin (Wang et al., 2023), Get3D (Gao et al., 2022), Zero123 (Liu et al., 2023d), SyncDreamer (Liu et al., 2023b), and LRM ((Hong et al., 2023)), etc.
@@
Instead of generating 3D through 2D image lifting, another research direction involves directly learning to produce 3D data, using supervision from either 3D shape ground truths or posed multi-view images. Notable methods include EG3D (Chan et al., 2022), Rodin (Wang et al., 2023), and Get3D (Gao et al., 2022), etc. These approaches represent 3D data with a continuous hybrid neural field architecture, typically consisting of a tri-plane and an MLP decoder. The tri-plane is a tensor used to factorize the dense 3D volume grid. It is built on three axis-aligned 2D planes: the XY, YZ, and XZ planes. The MLP decoder converts the tri-plane feature into a continuous value representing the scene, which could be occupancy, signed distance field (SDF), radiance field (Mildenhall et al., 2021), etc. Tri-plane is significantly more compact and computationally efficient than a full 3D tensor and conducive to generative architectures developed for 2D image synthesis. This has been a key factor in making high-quality direct 3D data generation possible.
@@ 2.1. Diffusion models
The success of 2D generation tools based on diffusion models, notably Stable Diffusion(Rombach et al., 2022b), Midjourney, and Dall-E, has significantly sparked interest in the development of 3D generation tools. There are two main streams for this task: the methods that lift 2D (generated) images into 3D models, and the methods that directly run diffusion on 3D data.
@@
3D diffusion models. Another line of research involves directly training diffusion models to generate 3D shapes. As the supervision comes directly from 3D shape ground truth or posed multi-view images, the generated results typically exhibit superior geometric quality compared to those from 2D diffusion-based methods. These methods can be categorized based on the type of 3D representations they employ, including: polygon meshes (Gao et al., 2022; Liu et al., 2023a), point clouds (Nichol et al., 2022; Zeng et al., 2022), explicit 3D grids holding occupancy or SDF values (Zheng et al., 2023; Liu et al., 2023c), or neural fields (Wang et al., 2023; Shue et al., 2023; Chen et al., 2023a; Müller et al., 2023; Jun and Nichol, 2023; Xu et al., 2023b; Erkoç et al., 2023; Chou et al., 2023). The hybrid neural field, which incorporates a tri-plane followed by a neural decoder, has been widely adopted in 3D diffusion models due to its computational efficiency. Rodin (Wang et al., 2023) first fits tri-plane NeRFs for a human upper body dataset, and then uses a two-stage coarse-to-fine diffusion model to generate the corresponding tri-planes. Similarly, NFD (Shue et al., 2023) trains a tri-plane diffusion model for 3D data parametrized via occupancy values. SSDNerf (Chen et al., 2023a) merge tri-plane fitting and generation into a single-stage pipeline. However, in practice, tri-plane diffusion is still challenging to train due to its high dimensionality and irregularity. Existing methods only demonstrated simple cases with small data varieties, i.e., Rodin for canonicalized human upper-body dataset (Wood et al., 2021), and NFD and SSDNeRF for single-category objects in ShapeNet (Chang et al., 2015). This paper follows this line of research but introduces a major change: we use an auto-encoder to compress the tri-plane into a highly compact latent tri-plane space for diffusion. We demonstrate that this approach significantly improves the stability, generalizability, and output quality of tri-plane diffusion.
@@
Generating 3D scenes presents a more substantial challenge than generating single objects. This is because scenes are geometrically more complex than individual objects, and they cannot be contained in a fixed spatial size. Object retrieval-based approaches assume there is a database of objects, and they arrange the retrieved objects to fill an empty scene, as seen in Diffuscene (Tang et al., 2023a) and Sceneformer (Wang et al., 2021), consequently, the synthesized scene can not contain novel elements that do not exist in the database. Text2Room (Höllein et al., 2023) is the first method that uses 2D diffusion model to build a 3D generation tool. It first generates color and depth frames using 2D diffusion models, and then shift camera positions to generate new frames, which are integrated into a global map. A similar approach for broader outdoor scenarios can be found in SceneScape (Fridman et al., 2023). To allow precise control over the contents generated in a scene, ControlRoom3D(Schult et al., 2023) and CTRL-ROOM(Fang et al., 2023) develop panorama-based room generation models that take 3D room layouts as input conditions. Citygen (Deng et al., 2023) represents city scenes using the height map proxy, leading to a 2.5D scene generation. Other approaches focus on generating scenes with high-quality visual appearances. Given a room scene mesh, MVDiffusion (Tang et al., 2023b) generates coherent multiview perspective images, which can be lifted to the 3D as the UV texture of the mesh. SceneDreamer (Chen et al., 2023c) leverages in-the-wild 2D images to construct large scenes with photo-realistic volume rendering effects. However, it still relies on high-quality 3D scene meshes as input, consequently, the dimensions of the generated scenes are bounded by the dimensions of the input meshes.
@@ 3.3. Compressing to latent tri-plane space
Although our raw tri-planes can reconstruct high-quality shapes, we found that generating such tri-planes is significantly difficult. Directly training diffusion models on such tri-planes leads to collapsed results, as shown in Fig. 4. We argue that there are mainly two reasons for this: 1) the raw tri-plane is highly redundant, and 2) the shape diversity in our scene block dataset is too large. Although previous works like Rodin (Wang et al., 2023) and NFD (Shue et al., 2023) have proven the feasibility of diffusion on raw tri-planes, they only work on datasets with much smaller varieties, i.e., Rodin for canonicalized human upper bodies, and NFD for single-category objects from ShapeNet (Chang et al., 2015). When we attempted to retrain NFD on our scene blocks, it also failed to produce meaningful shapes, as shown in Fig. 4.
We need to find a feature representation for 3D shapes that is compact, easy to train diffusion models, memory and computationally efficient, and capable of generalizing to large shape variations. In the 2D scenario, Stable Diffusion (Rombach et al., 2022a) compresses raw images into a latent 2D feature space for diffusion. This approach results in a more robust model that generates higher-quality images. Inspired by Stable Diffusion, we train an auto-encoder to compress raw tri-planes into a latent tri-plane space with reduced resolution and feature channels. Precisely, given an raw tri-plane $x\in{\mathbb{R}^{3\times{N^{2}}\times{C}}}$, the encoder $\mathcal{E}$ encodes $x$ into a latent representation $z=\mathcal{E}(x)$, and the decoder $\mathcal{D}$ reconstructs the raw tri-plane from the latent, giving $\hat{x}=\mathcal{D}(z)=\mathcal{D}(\mathcal{E}(x))$.
@@ 3.4. Latent Triplane Diffusion
3D aware denoising U-Net. The neural backbone $\Psi(\cdot)$ of our model is realized as a time-conditional U-Net. The advantage of tri-planes is that we can treat them as 2D tensors and therefore apply efficient 2D convolutions. However, naively running convolution on tri-planes does not produce satisfactory results, as the 3D relationships among the plane features are ignored. $\Psi(\cdot)$ needs to incorporate operations that can account for the cross-plane feature relationships. To address this, Rodin (Wang et al., 2023) introduces a 3D-aware convolution, which employs max-pooling and concatenation to associate features between planes based on their 3D correlations. However, the simple max-pooling may lose valuable information. In this work, we build $\Psi(\cdot)$ by leveraging the more powerful transformer to achieve cross-plane communication. The overall architecture of $\Psi(\cdot)$ is shown in Fig. 5. This architecture enables effective 3D-aware feature learning.
@@ 4. Experimental Results
Datasets. We test our algorithm on three different types of scenes: room, city, and village. Room scene data is obtained from 3DFront (Fu et al., 2021a) and 3D-FUTURE (Fu et al., 2021b), which contains 18,968 indoor scenes with 34 classes of indoor objects. We obtain 57K random crops from 3DFront, with each block size set to $3.2^{3}$ cubic meters. We filtered out empty rooms and rooms with less than 5 objects and finally got 9123 rooms. For simplicity, we regroup the objects in 3DFUTURE based on their similarities into 9 classes: ”floor”, ”wall”, ”chair”, ”cabinet”, ”sofa”, ”table”, ”lighting”, ”bed”, and ”stool”. The city and village scenes are designed by artists, from each, we obtain 10K blocks. The block sizes are set to $12^{3}$ and $15^{3}$ cubic meters, respectively. The layout labels for the village are ”pine”, ”cypress tree”, ”ground”, and ”houses”, and for the city, they are ”road”, ”tree”, ”solar panels”, ”cars”, ”houses”. Note that all the blocks are cropped at random, and the testing blocks are never exposed to the model during training.
Our method is implemented using Pytorch and trained on Nvidia V100 GPU. For the 3DFront dataset with 57K cropped blocks, raw tri-plane fitting, auto-encoder training, and diffusion training take 4750, 768, and 384 GPU hours, respectively. Running a single tri-plane extrapolation under layout conditions costs 6 minutes. With the large scene generation strategy described in Sec. 3.7, producing the large indoor scene in Fig. 7 takes around 3 hours.
@@ 4.4. Ablation study
Figure 13. Left: latent tri-plane extrapolation result, rights: after applying non-rigid registration.
@@ 5. Conclusion and Discussion
Limitations. The current implementation of BlockFusion faces several limitations. Our method may fail to generate very fine geometric details in the scene, such as the legs of a chair. This issue primarily stems from the limited resolution used for the tri-planes. A possible solution is to adopt tri-plane super-resolution. Moreover, the bounding box condition can only control the approximate placement of objects, not their orientations. We believe that precise orientation control could be achieved by training diffusion conditioning on both the bounding box map and an object orientation map. This orientation map can also be easily obtained from user instructions. Lastly, while we have demonstrated textured mesh results on small scenes, the task of generating globally consistent textures for large scene meshes is both a challenging and intriguing future endeavor.

### 09 Apr 2024 v3

@@ Abstract.
ccs: Computing methodologies Shape modelingccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Generating large amount of high-quality 3D contents is key for many practical applications, including video-games, film-making, augmented and virtual reality (AR/VR). The increasing demand for high-quality digital contents has made 3D generation a significant topic of research. Recently, in the 2D domain, denoise diffusion models (Sohl-Dickstein et al., 2015), have demonstrated remarkable results in image synthesis and beyond, leading to the development of production-ready 2D generation tools, such as Stable Diffusion (Rombach et al., 2022a), Midjourney, and Dall-E (Ramesh et al., 2021). The success in 2D domain has significantly sparked interest in the development of 3D generation tools. A multitude of researches on 3D generation have been published recently, most notable works include DreamFusion (Poole et al., 2022), Rodin (Wang et al., 2023b), Get3D (Gao et al., 2022), Zero123 (Liu et al., 2023d), SyncDreamer (Liu et al., 2023b), and LRM ((Hong et al., 2023)), etc.
@@
Instead of generating 3D through 2D image lifting, another research direction involves directly learning to produce 3D data, using supervision from either 3D shape ground truths or posed multi-view images. Notable methods include EG3D (Chan et al., 2022), Rodin (Wang et al., 2023b), and Get3D (Gao et al., 2022), etc. These approaches represent 3D data with a continuous hybrid neural field architecture, typically consisting of a tri-plane and an MLP decoder. The tri-plane, originally introduced in EG3D (Chan et al., 2022) and (Peng et al., 2020), is a tensor used to factorize the dense 3D volume grid. It is built on three axis-aligned 2D planes: the XY, YZ, and XZ planes. The MLP decoder converts the tri-plane feature into a continuous value representing the scene, which could be occupancy, signed distance field (SDF), radiance field (Mildenhall et al., 2021), etc. Tri-plane is significantly more compact and computationally efficient than a full 3D tensor and conducive to generative architectures developed for 2D image synthesis. This has been a key factor in making high-quality direct 3D data generation possible.
@@ 2.1. Diffusion models
The success of 2D generation tools based on diffusion models, notably Stable Diffusion(Rombach et al., 2022b), Midjourney, and Dall-E, has significantly sparked interest in the development of 3D generation tools. There are two main streams for this task: the methods that lift 2D (generated) images into 3D models, and the methods that directly run diffusion on 3D data. A thorough review of diffusion models for visual computing is available in (Po et al., 2023).
@@
3D diffusion models. Another line of research involves directly training diffusion models to generate 3D shapes. As the supervision comes directly from 3D shape ground truth or posed multi-view images, the generated results typically exhibit superior geometric quality compared to those from 2D diffusion-based methods. These methods can be categorized based on the type of 3D representations they employ, including: polygon meshes (Gao et al., 2022; Liu et al., 2023a), point clouds (Nichol et al., 2022; Zeng et al., 2022), explicit 3D grids holding occupancy or SDF values (Zheng et al., 2023; Liu et al., 2023c), or neural fields (Wang et al., 2023b; Shue et al., 2023; Chen et al., 2023a; Müller et al., 2023; Jun and Nichol, 2023; Xu et al., 2023b; Erkoç et al., 2023; Chou et al., 2023). The hybrid neural field, which incorporates a tri-plane followed by a neural decoder, has been widely adopted in 3D diffusion models due to its computational efficiency. Rodin (Wang et al., 2023b) first fits tri-plane NeRFs for a human upper body dataset, and then uses a two-stage coarse-to-fine diffusion model to generate the corresponding tri-planes. Similarly, NFD (Shue et al., 2023) trains a tri-plane diffusion model for 3D data parametrized via occupancy values. SSDNerf (Chen et al., 2023a) merge tri-plane fitting and generation into a single-stage pipeline. However, in practice, tri-plane diffusion is still challenging to train due to its high dimensionality and irregularity. Existing methods only demonstrated simple cases with small data varieties, i.e., Rodin for canonicalized human upper-body dataset (Wood et al., 2021), and NFD and SSDNeRF for single-category objects in ShapeNet (Chang et al., 2015). This paper follows this line of research but introduces a major change: we use an auto-encoder to compress the tri-plane into a highly compact latent tri-plane space for diffusion. We demonstrate that this approach significantly improves the stability, generalizability, and output quality of tri-plane diffusion.
@@
Generating 3D scenes presents a more substantial challenge than generating single objects. This is because scenes are geometrically more complex than individual objects, and they cannot be contained in a fixed spatial size. Object retrieval-based approaches assume there is a database of objects, and they arrange the retrieved objects to fill an empty scene, as seen in Diffuscene (Tang et al., 2023a) and Sceneformer (Wang et al., 2021), consequently, the synthesized scene can not contain novel elements that do not exist in the database. Text2Room (Höllein et al., 2023) is the first method that uses 2D diffusion model to build a 3D generation tool. It first generates color and depth frames using 2D diffusion models, and then shift camera positions to generate new frames, which are integrated into a global map. A similar approach for indoor scenarios can be found in SceneScape (Fridman et al., 2023). To allow precise control over the contents generated in a scene, ControlRoom3D(Schult et al., 2023) and CTRL-ROOM(Fang et al., 2023) develop panorama-based room generation models that take 3D room layouts as input conditions. CC3D (Bahmani et al., 2023) utilizes a 3D layout to guide the SDS (Poole et al., 2022) process. Citygen (Deng et al., 2023) represents city scenes using the height map proxy, leading to a 2.5D scene generation. Other approaches, such as (Tang et al., 2023b; Chen et al., 2023c; Wang et al., 2023a) focus on generating scenes with high-quality visual appearances. Given a room scene mesh, MVDiffusion (Tang et al., 2023b) generates coherent multiview perspective images, which can be lifted to the 3D as the UV texture of the mesh.
SceneDreamer (Chen et al., 2023c) use in-the-wild 2D image collections to learn a feed-forward generative model for unbounded 3D scenes. Semantic height maps are required while inference.
@@ 3.3. Compressing to latent tri-plane space
Although our raw tri-planes can reconstruct high-quality shapes, we found that generating such tri-planes is significantly difficult. Directly training diffusion models on such tri-planes leads to collapsed results, as shown in Fig. 4. We argue that there are mainly two reasons for this: 1) the raw tri-plane is highly redundant, and 2) the shape diversity in our scene block dataset is too large. Although previous works like Rodin (Wang et al., 2023b) and NFD (Shue et al., 2023) have proven the feasibility of diffusion on raw tri-planes, they only work on datasets with much smaller varieties, i.e., Rodin for canonicalized human upper bodies, and NFD for single-category objects from ShapeNet (Chang et al., 2015). When we attempted to retrain NFD on our scene blocks, it also failed to produce meaningful shapes, as shown in Fig. 4.
We need to find a feature representation for 3D shapes that is compact, easy to train diffusion models, memory and computationally efficient, and capable of generalizing to large shape variations. In the 2D scenario, Stable Diffusion (Rombach et al., 2022a) compresses raw images into a latent 2D feature space for diffusion. This approach results in a more robust model that generates higher-quality images. Inspired by Stable Diffusion, we train an auto-encoder to compress raw tri-planes into a latent tri-plane space with reduced resolution and feature channels. Precisely, given a raw tri-plane $x\in{\mathbb{R}^{3\times{N^{2}}\times{C}}}$, the encoder $\mathcal{E}$ encodes $x$ into a latent representation $z=\mathcal{E}(x)$, and the decoder $\mathcal{D}$ reconstructs the raw tri-plane from the latent, giving $\hat{x}=\mathcal{D}(z)=\mathcal{D}(\mathcal{E}(x))$.
@@ 3.4. Latent Triplane Diffusion
3D aware denoising U-Net. The neural backbone $\Psi(\cdot)$ of our model is realized as a time-conditional U-Net. The advantage of tri-planes is that we can treat them as 2D tensors and therefore apply efficient 2D convolutions. However, naively running convolution on tri-planes does not produce satisfactory results, as the 3D relationships among the plane features are ignored. $\Psi(\cdot)$ needs to incorporate operations that can account for the cross-plane feature relationships. To address this, Rodin (Wang et al., 2023b) introduces a 3D-aware convolution, which employs max-pooling and concatenation to associate features between planes based on their 3D correlations. However, the simple max-pooling max-pooling selects the largest one and discards the rest, inevitably causing information loss. In this work, we build $\Psi(\cdot)$ by leveraging the more powerful transformer to achieve cross-plane communication. The overall architecture of $\Psi(\cdot)$ is shown in Fig. 5. This architecture enables effective 3D-aware feature learning.
@@ 4. Experimental Results
Datasets. We test our algorithm on three different types of scenes: room, city, and village. Room scene data is obtained from 3DFront (Fu et al., 2021a) and 3D-FUTURE (Fu et al., 2021b), which contains 18,968 indoor scenes with 34 classes of indoor objects. We obtain 57K random crops from 3DFront, with each block size set to $3.2^{3}$ cubic meters. We filtered out empty rooms and rooms with less than 5 objects and finally got 9123 rooms. For simplicity, we regroup the objects in 3DFUTURE based on their similarities into 9 classes: ”floor”, ”wall”, ”chair”, ”cabinet”, ”sofa”, ”table”, ”lighting”, ”bed”, and ”stool”. The city and village scenes are designed by artists, from each, we obtain 10K blocks. The block sizes are set to $12^{3}$ and $15^{3}$ cubic meters, respectively. The layout labels for the village are ”pine”, ”cypress tree”, ”ground”, and ”houses”, and for the city, they are ”road”, ”tree”, ”solar panels”, ”cars”, ”houses”. Note that all the blocks are cropped at random, and the testing blocks are never exposed to the model during training. During inference, the input semantic layout is created using an easy-to-use GUI, where user can place bounding boxes to indicate objects such as ”car” and ”tree”, or draw lines/contours to indicate continuous areas such as ”wall” and ”road”.
Our method is implemented using Pytorch and trained on Nvidia V100 GPU. For the 3D Front dataset with 57K cropped blocks, raw tri-plane fitting, auto-encoder training, and diffusion training take 4750, 768, and 384 GPU hours, respectively. VAE and diffusion training require 8 GPUs. Tri-plane fitting and BlockFusion inference can run on one GPU. Running a single tri-plane extrapolation under layout conditions costs 6 minutes. With the large scene generation strategy described in Sec. 3.7, producing the large indoor scene in Fig. 7 takes around 3 hours.
@@ 4.4. Ablation study
Figure 13. Left: latent tri-plane extrapolation result (the seams are more visible when zoomed in), rights: after applying non-rigid registration.
@@ 5. Conclusion and Discussion
Advantage over procedural generation.
Procedural Content Generation (PCG) is a complicated system that heavily depends on expertise. It requires carefully crafted rules to generate meaningful 3D scene. These rules must be re-programmed when transitioning between different scene styles. In contrast, BlockFusion is rule-programming-free and easy to use. It learns the distribution of scenes directly from data, and the generation process is fully automated.
Limitations.
The current implementation of BlockFusion faces several limitations. Our method may fail to generate very fine geometric details in the scene, such as the legs of a chair. This issue primarily stems from the limited resolution used for the tri-planes. A possible solution is to adopt tri-plane super-resolution. Moreover, the bounding box condition can only control the approximate placement of objects, not their orientations. We believe that precise orientation control could be achieved by training diffusion conditioning on both the bounding box map and an object orientation map. This orientation map can also be easily obtained from user instructions. Lastly, while we have demonstrated textured mesh results on small scenes, the task of generating globally consistent textures for large scene meshes is both a challenging and intriguing future endeavor.